jgate:

Create a unique parabola in the pattern f(x) = (x − a)(x − b). Describe the direction of the parabola and determine the y-intercept and zeros.

WickedHeart:

Hello what would you say the answer would be?

surjithayer:

\[f(x)=(x-a)(x-b)=x^2-ax-bx+ab\] \[f(x)=x^2-(a+b)x+ab\] \[x^2>0\] so parabola is open above. when x=0,f(x)=ab y-intercept is (0,ab) zeros are x=a,b